OpenStudy (anonymous):

5th grade work PLZZ HELPP i never learned these before (summer packet and i have only one day to finshi it until school starts tommmomroow and any one know some math free tutor website Q: 4+(-12)= -12+(-14)= 10+(-1)= -2+(-1)= -50+(-1)= 8+(-4)= -9+7= its adding integers

OpenStudy (anonymous):

Hint: Adding a negative number is the same as subtracting that number. Example: 5+(-1) is the same as 5-1 which is 4.

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

You could think of positive and negative numbers in terms of money. Positive is what you have and negative is what you owe. So if you have $5 and owe $1(to pay back), you have $4. 5 +(-1) = 4 If you owe $2 and owe $1, you owe $3. -2 + (-1) = -3 If you owe $9 and have $7 (to pay back) , you owe $2. -9 + 7 = -2 Does this help it to make sense?

OpenStudy (anonymous):

Or remember this: + + = + + - = - - + = - - - = + So if you subtract a negative number you actually add it, etc.

OpenStudy (anonymous):

wel year but how u get the answer of it catherine and traxter ohh but i dont stilll get how to get the answer tho

OpenStudy (anonymous):

Traxter gave you some rules to memorize; I tried to give you everyday examples that you would be familiar with.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

If you owe me $5, and you owe Traxter $3, how much do you owe?

OpenStudy (anonymous):

6$

OpenStudy (anonymous):

Not quite, try again.

OpenStudy (anonymous):

Oops... try again. (You owe $5 and you owe $3)

OpenStudy (anonymous):

my mistake is 8$ my mind stupid and have million thing going on

OpenStudy (anonymous):

You just did it! -5 + (-3) = -8

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh yeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeesssssssssssssssssssssssssssssssss i ll try to do it my homework and and ill tell u the answer if is right

OpenStudy (anonymous):

:-D

OpenStudy (anonymous):

:))